OpenStudy (anonymous):

lim t → infinity ((SQROOT of t) + t^2)/(8t − t^2)

myininaya (myininaya):

Since the highest exponent in the denominator is 2. We will divide both top and bottom by t^2 You should know that \[\lim_{x \rightarrow \infty}1/x^p=0, p \ge 1\]

myininaya (myininaya):

I assume you got your one question since you closed it?

OpenStudy (anonymous):

no i didn't :/

myininaya (myininaya):

Ok. Well how about you focus on it before opening more questions. I'm going to leave soon and I would like to finish that problem with you before I go.

OpenStudy (anonymous):

ok but i need this one too

OpenStudy (anonymous):

can we do both please